How many MPH (to the nearest 100th) is a car traveling if the wheels have a diameter of 3 feet and they are turning 600 times a minute?
Answer by mananth(16946)

How many MPH (to the nearest 100th) is a car traveling if the wheels have a diameter of 3 feet and they are turning 600 times a minute?
diameter of wheel = 3 feet
circumference of wheel = pi*D
= pi*3
=9.42
one revolution it covers a distance of 9.42 feet
600 revolutions it covers a distance of 600 * 9.42
= 5654.86 feet - this distance the car covers in 1 minute
so in 60 minutes it covers a distance of 5654.86*60
=339,292 feet / hour
= 339,292 / 5280 = 64.26 mph (5280 feet = 1 mile)
Ananth